Higher education. The table lists thje average costs of attending a four year college in the U.S. during the 2005-2006 school year

College Sector Tuition and fees room and board
1-year public $5941 $6636

1-year private $21,235 $7,791

Russ's parents invested money to a savings account earning an average of 4.5 percent interest, compounded monthly.
After 15 years they have exactly the right amount to cover the tuition, fees, room, and board for Russ's first year at a public college. What was their initial investment? Round your answer to the nearest dollar.
What I tried:
One year in public is 5941+6636=12577
I plugged into the compound interest formula: A=P(1+(r/n))^nt which is
12577=P(1+0.045/12)^12(5)
12577=P(1.00375)^60
12577=1.2517958205P
10031 is about P
I think I did something wrong, as the answer doesn't make sense.
Thanks

Answer by addingup(3677)

5941+6636 = 12577
P(1+(0.045/12))^12*15 = 12577
P(1.00375)^180 = 12577
P*1.9616 = 12577
P = 6411.60
.
NOTE: In your calculations you used 5 years but in the problem you posted you say 15 years, so I used 15 years. By the way, there's nothing wrong with your calculations.
.
Here it is with 5 years:
P(1+(0.045/12))^12*5
P(1.00375)^60
P*1.2518 = 12,557
P = 10,031.16 (same answer you got, this is how much they initially saved)